The equation of motion of a particle moving along the x-axis is given by: d²x/dt² + 4dx/dt + 8x = 20cos2t. Calculate the amplitude, period and frequency of the oscillation after a long time has elapsed
1
Expert's answer
2020-02-24T10:04:51-0500

The equation of motion of a particle moving along the x-axis is given by


"\frac{d²x}{dt²} + 4\frac{dx}{dt} + 8x = 20\cos2t."

The solution


"x=A\cos2t+B\sin2t."


We obtain


"-4A\cos2t-4B\sin2t - 8A\sin2t+8B\cos2t""+ 8A\cos2t+8B\sin2t = 20\cos2t."

Hence


"4A+8B=20,\quad -8A+4B=0.""B=2A, \; A=1.""x=\cos2t+2\sin2t=\sqrt{5}\cos(2t+\arctan2)."

Therefore:

amplitude "a=\sqrt{5};"

period "T=\frac{2\pi}{\omega}=\pi;"

frequency "f=1/T=1/\pi."